a speaker from brand A can generate a sound level of 60dB what is the sound level of a speaker from brand Barber if it can generate sound that has 300 times higher intensity than that of brand A
Answer #1

To me it seems a very straight forward question.

As intensity of A is 60 db and intensity of B is 300 times greater than A,

Intensity of A in watt/m2 = 0.000001 watt/m2

As from the formula I(db) = 10Log I/Io Here I = intensity, and Io = reference intensity = 10^-12 watt/m2

So, intensity of B = 300x0.000001 = 0.003 watt/m2

So, from the formula again I(db) of B = 10 Log(0.003/10^-12)

I(db)B = 10*8.47 = 84.7 db Ans.